I am six years older than my brother. There was a time when I was twice as old as he was. What was
his age when I was twice as old?

Answer by ewatrrr(24785)

HI,
*Note: Your have been and will "always" be 6 years older than your brother
Let x be your brothers age and (x + 6) be your age at the time you were twice his age.
.
(x+6)= 2x
x + 6 = 2x
6 = x
.
your brother: 6 years old at the time you were twice as old.
.
Check yur answer
2*6 = (6 + 6 )